When turning at 1500 rev/min, a certain generator produces 100.0v. What must be its frequency in rev/min if it is to produce 120.0v

1
Expert's answer
2022-04-05T19:19:23-0400
"{\cal E}=2\pi n NBA"

"\frac{\cal E_1}{\cal E_2}=\frac{n_1}{n_2}"

"n_2=\frac{\cal E_2}{\cal E_1}n_1=\frac{120}{100}*1500=1800\:\rm rev/min"